# On the complete separation of unique $\ell_{1}$ spreading models and the Lebesgue property of Banach spaces

### v1

@@ 1. Introduction
Let us explain the main properties of $X_{\mathcal{D}}$. Denote by $\mathcal{D}$ the binary tree $\cup_{n=0}^{\infty}\{0,1\}^{n}$. Following standard convention, $\{0,1\}^{0}$ is a singleton containing the empty sequence. For $n\in\mathbb{N}\cup\{0\}$ and $\lambda\in\{0,1\}^{n}$ we write $|\lambda|=n$ and say $\lambda$ has height $n$. We consider $\mathcal{D}$ with the usual partial order with minimum the empty sequence and, otherwise, $\lambda\leq\mu$ means $|\lambda|\leq|\mu|$ and $\lambda(i)=\mu(i)$, for $1\leq i\leq|\lambda|$. The characterization of the Lebesgue property in terms of Haar-$\ell_{1}^{+}$ sequences from [11] easily yields the following (see Remark 2.4): a Banach space $X$ has the Lebesgue property if and only if for every collection $(x_{\lambda})_{\lambda\in\mathcal{D}}$ of norm-one vectors in $X$ there exist $\theta>0$ and, for every $n\in\mathbb{N}$, a collection $(\mu_{\lambda})_{\lambda\in\{0,1\}^{n}}$ such that, for $\lambda\in\{0,1\}^{n}$, $\mu_{\lambda}\geq\lambda$ and
$\frac{1}{2^{n}}\Big{\|}\sum_{\lambda\in\{0,1\}^{n}}x_{\mu_{\lambda}}\Big{\|}\geq\theta.$
In the space $X_{\mathcal{D}}$, every infinite-dimensional subspace contains a lexicographically block collection $(x_{\lambda})_{\lambda\in\{0,1\}^{n}}$ of norm-one vectors such that for any $n\in\mathbb{N}$ and collection $(\mu_{\lambda})_{\lambda\in\{0,1\}^{n}}$ in $\mathcal{D}$ such that $\mu_{\lambda}\geq\lambda$, $\lambda\in\{0,1\}^{n}$,
(1) $\max_{\lambda\in\{0,1\}^{n}}|a_{\lambda}|\leq\Big{\|}\sum_{\lambda\in\mathcal{D}^{n}}a_{\lambda}x_{\mu_{\lambda}}\Big{\|}\leq 3\max_{\lambda\in\mathcal{D}^{n}}|a_{\lambda}|,$
for any scalars $(a_{\lambda})_{\lambda\in\mathcal{D}^{n}}$. Therefore, $X_{\mathcal{D}}$ fails the Lebesgue property. At the same time, every normalized weakly null sequence in $X_{\mathcal{D}}$ has a subsequence generating a spreading model that is $102$-equivalent to the unit vector basis of $\ell_{1}$. In particular, it from follows James’s characterization of reflexivity for spaces with an unconditional basis that $X_{\mathcal{D}}$ is reflexive.
@@ 2. Preliminary information
    For every collection $(x_{\lambda})_{\lambda\in\mathcal{D}}$ of norm-one vectors in $X$, there exists a constant $\theta>0$ such that, for every $n\in\mathbb{N}$ and for every $\lambda\in\{0,1\}^{n}$, we have that
    $\frac{1}{2^{n}}\left\|\sum_{\lambda\in\{0,1\}^{n}}x_{\mu_{\lambda}}\right\|\geq\theta$
    for some collection of nodes $(\mu_{\lambda})_{\lambda\in\{0,1\}^{n}}$ with $\mu_{\lambda}\geq\lambda$.
@@ 3. Construction of $X_{\mathcal{D}}$ and uniformly unique $\ell_{1}$ spreading model
where $w(f)=m_{j_{1}}\cdots m_{j_{l}}$, $\vec{w}(f)=(m_{j_{1}},\ldots,m_{j_{\ell}})$, and $m(f)=\max_{1\leq z\leq l}m_{j_{z}}$ for $f\in W_{1}\setminus W_{0}$. We next define the following notion of proximity with respect to $\overline{\mathcal{D}}$ for functionals in $W_{1}$.
@@
    If $f_{1}<\cdots<f_{d}$ is a $(n_{j_{1}},\ldots,n_{j_{\ell}},\lambda)$-packed sequence in $W$ and $g_{1}<\cdots<g_{d}$ are in $c_{00}$ such that, for $1\leq i\leq d$ and $s\in\mathrm{supp}(g_{i})$, we have $|g_{i}(e_{s})|=|f_{i}(e_{s})|$ then $g_{1}$,…,$g_{d}$ are also in $W$ and they are $(n_{j_{1}},\ldots,n_{j_{\ell}},\lambda)$-tightly packed.
@@ 4. Auxiliary Banach spaces and $c_{0}$-type estimate
[…]
@@
[…]
@@ 5. Rapidly increasing sequences and basic inequality
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ 1. Introduction
Let us explain the main properties of $X_{\mathcal{D}}$. Denote by $\mathcal{D}$ the binary tree $\cup_{n=0}^{\infty}\{0,1\}^{n}$. Following standard convention, $\{0,1\}^{0}$ is a singleton containing the empty sequence. For $n\in\mathbb{N}\cup\{0\}$ and $\lambda\in\{0,1\}^{n}$ we write $|\lambda|=n$ and say $\lambda$ has height $n$. We consider $\mathcal{D}$ with the usual partial order with minimum the empty sequence and, otherwise, $\lambda\leq\mu$ means $|\lambda|\leq|\mu|$ and $\lambda(i)=\mu(i)$, for $1\leq i\leq|\lambda|$. The characterization of the Lebesgue property in terms of Haar-$\ell_{1}^{+}$ sequences from [11] easily yields the following (see Remark 2.4): a Banach space $X$ has the Lebesgue property if and only if, for every collection $(x_{\lambda})_{\lambda\in\mathcal{D}}$ of norm-one vectors in $X$, there exists a constant $\theta>0$ such that, for every $n\in\mathbb{N}$, there exists a subset $V\subset\{0,1\}^{n}$ such that
$\theta\leq\frac{1}{2^{n}}\left\|\sum_{\lambda\in V}x_{\mu_{\lambda}}\right\|$
for some choice of nodes $(\mu_{\lambda})_{\lambda\in V}$ with $\mu_{\lambda}\geq\lambda$. In the space $X_{\mathcal{D}}$, every infinite-dimensional subspace contains a lexicographically block collection $(x_{\lambda})_{{\lambda\in\mathcal{D}}}$ of norm-one vectors such that for any $n\in\mathbb{N}$ and collection $(\mu_{\lambda})_{\lambda\in\{0,1\}^{n}}$ in $\mathcal{D}$ such that $\mu_{\lambda}\geq\lambda$, $\lambda\in\{0,1\}^{n}$,
(1) $\max_{\lambda\in\{0,1\}^{n}}|a_{\lambda}|\leq\Big{\|}\sum_{{\lambda\in\{0,1\}^{n}}}a_{\lambda}x_{\mu_{\lambda}}\Big{\|}\leq 3\max_{{\lambda\in\{0,1\}^{n}}}|a_{\lambda}|,$
for any scalars $(a_{\lambda})_{{\lambda\in\{0,1\}^{n}}}$. Therefore, $X_{\mathcal{D}}$ fails the Lebesgue property. At the same time, every normalized weakly null sequence in $X_{\mathcal{D}}$ has a subsequence generating a spreading model that is $102$-equivalent to the unit vector basis of $\ell_{1}$. In particular, it follows from James’s characterization of reflexivity for spaces with an unconditional basis that $X_{\mathcal{D}}$ is reflexive.
@@ 2. Preliminary information
    For every collection $(x_{\lambda})_{\lambda\in\mathcal{D}}$ of norm-one vectors in $X$, there exists a constant $\theta>0$ such that, for every $n\in\mathbb{N}$, there exists a subset $V\subset\{0,1\}^{n}$ such that
    $\theta\leq\frac{1}{2^{n}}\left\|\sum_{\lambda\in V}x_{\mu_{\lambda}}\right\|$
    for some choice of nodes $(\mu_{\lambda})_{\lambda\in V}$ with $\mu_{\lambda}\geq\lambda$.
@@ 3. Construction of $X_{\mathcal{D}}$ and uniformly unique $\ell_{1}$ spreading model
where $w(f)=m_{j_{1}}\cdots m_{j_{l}}$, $\vec{w}(f)=(m_{j_{1}},\ldots,m_{j_{\ell}})$, and $m(f)=\max_{1\leq z\leq l}m_{j_{z}}$ for $f\in W_{1}\setminus W_{0}$. For $f$ in $W_{1}\setminus W_{0}$, $\vec{w}(f)$ may not be uniquely defined, and thus neither are $w(f)$ and $m(f)$. The notation $\vec{w}(f)=(m_{j_{1}},\ldots,m_{j_{\ell}})$ means that $f$ admits a representation as above, and it is fixed while working with $f$. This will also apply to $f\in W_{k}$, $k>1$. This is common in this type of construction and does not pose a problem. In fact, insisting on uniquely defining the weight may interfere with constructing functionals that are used to prove the properties of the space. We next define the following notion of proximity with respect to $\overline{\mathcal{D}}$ for functionals in $W_{1}$.
@@
    If $f_{1}<\cdots<f_{d}$ is a $(n_{j_{1}},\ldots,n_{j_{\ell}},\lambda)$-packed sequence in $W$ and $g_{1}<\cdots<g_{d}$ are in $c_{00}$ such that, for $1\leq i\leq d$ and $s\in\mathrm{supp}(g_{i})$, we have $|g_{i}(e_{s})|=|f_{i}(e_{s})|$ then $g_{1}$,…,$g_{d}$ are also in $W$ and they are $(n_{j_{1}},\ldots,n_{j_{\ell}},\lambda)$-packed.
@@ 4. Auxiliary Banach spaces and $c_{0}$-type estimate
Let $f=\frac{1}{m_{j_{1}}\cdots m_{j_{l}}}\sum_{i=1}^{d}f_{i}\in W^{\text{aux},\delta}$. Then, $\{i\mid f_{i}\notin W^{\text{aux},\delta}_{0}\}=E_{0}\cup E_{1}\cup\ldots\cup E_{l}$ where $\mathrm{card}(E_{0})\leq 1$ and, for each $1\leq z\leq l$, $(f_{i})_{i\in E_{z}}$ is $S_{n_{j_{1}}+\cdots+n_{j_{z}}}\ast\mathcal{A}_{3}$-admissible and $\max_{i\in E_{z}}d_{T}(m(f_{i}),\lambda_{z})\leq\delta$ for some $\lambda_{z}\in\overline{\mathcal{D}}$.
@@
for each $1\leq i\leq 2^{N}$. It follows that $\left|f\left(\sum_{i\in\Omega}a_{i}x_{i}\right)\right|\leq\left(1+\frac{4}{\sqrt{w(f)}}\right)\max_{i\in\Omega}|a_{i}|$ for every subset $\Omega\subset\{1,\ldots,2^{N}\}$, for every choice $(a_{i})_{i\in\Omega}$ of scalars, and for every auxiliary weighted functional $f=\frac{1}{m_{j_{1}}\cdots m_{j_{l}}}\sum_{q=1}^{d}f_{q}\in W^{\text{aux},\delta}$ where $\delta<2^{-(N+1)}$.
@@ 5. Rapidly increasing sequences and basic inequality
Fix $\delta>0$ and assume that $f_{1}<\ldots<f_{d}\in W$ with $2^{-\max\text{supp}(f_{1})}\leq\frac{\delta}{2}$ are $(n_{j_{1}},\ldots,n_{j_{l}},\lambda)$-packed for some $\lambda\in\overline{\mathcal{D}}$. Next, suppose that there exist sets of functionals $\mathcal{F}_{1},\ldots,\mathcal{F}_{d}$ such that, for each $1\leq q\leq d$, $\mathcal{F}_{q}=\{g^{q}_{1},g^{q}_{2},g^{q}_{3}\}$ with $g^{q}_{r}\in\{0\}\cup W^{\text{aux},\delta}$, and:
@@
    $\cup_{r=1}^{3}\text{supp}(g^{q}_{r})\subset\text{supp}(f_{q})$ and $\text{supp}(g^{q}_{1})<\text{supp}(g^{q}_{2})<\text{supp}(g^{q}_{3})$,
@@
Let $(x_{i})_{i\in I}$ be a $(C,(j_{i})_{i\in I})$-RIS with $2^{-\min\text{supp}(x_{\min(I)})}\leq\delta$. Then, for every functional $f=\frac{1}{m_{j_{1}}\cdots m_{j_{l}}}\sum_{q=1}^{d}f_{q}\in W$, there exist functionals $h\in\{0\}\cup W^{\text{aux},\delta}_{0}$ and $g\in W^{\text{aux},\delta}$ with $\vec{w}(g)=\vec{w}(f)$ such that
@@
for every choice $(a_{i})_{i\in I}$ of scalars where $t_{i}=\max\text{supp}(x_{i})$ for each $i\in I$.